# Do spectral cues matter in contrast-based graph self-supervised learning?

### v1

@@ Appendix B Dataset and training configuration
Datasets. The node classification datasets used in this paper include the Cora, CiteSeer, and PubMed citation networks [13], as well as the Photo and Computers co-purchase networks [27]. Additionally, we use the Coauthor-CS and Coauthor-Phy co-author relationship networks. The statistics of node-level datasets are present in Table 4. The graph classification datasets include: The MUTAG dataset, which features seven types of graphs derived from 188 mutagenic compounds; the NCI1 dataset, which contains compounds tested for their ability to inhibit human tumor cell growth; the PROTEINS dataset, where nodes correspond to secondary structure elements connected if they are adjacent in 3D space; and the IMDB-BINARY and IMDB-MULTI movie collaboration datasets, where graphs depict interactions among actors and actresses, with edges denoting their collaborations in films. These movie graphs are labeled according to their genres. The statistics of graph-level datasets are present in Table 5. All datasets can be accessed through PyG library ***https://pytorch-geometric.readthedocs.io/en/latest/modules/datasets.html. All experiments are conducted using 8 NVIDIA A100 GPU.
@@
Training configuration. For each CG-SSL framework, we implement it based on [41]  https://github.com/PyGCL/PyGCL. We use the following hyperparameters: the learning rate is set to $5\times 10^{-4}$, and the node hidden size is set to $512$, the number of GCN encoder layer is set $\in\{1,2\}$. For all node classification datasets, training epochs are set $\in\{50,100,150,200,400,1000\}$, and for all graph classification datasets, training epochs are set $\in\{20,40,...,200\}$. To achieve performance closer to the global optimum, we use randomized search to determine the optimal probability of edge perturbation and SPAN perturbation ratio. For Cora and CiteSeer  the search is conducted one hundred times, and for all other datasets, it is conducted twenty times. For all graph classification datasets, the batch size is set to $128$.
Appendix C Preliminaries of Graph Spectrum and SPAN
@@
$\max_{\boldsymbol{\mathcal{T}}_{1},\boldsymbol{\mathcal{T}}_{2}\in\mathcal{S}}\left\|\operatorname{eig}\left(\mathbf{L}_{1}\right)-\operatorname{eig}\left(\mathbf{L}_{2}\right)\right\|_{2}^{2}$ (7)
@@
Appendix D Object function of GCL framework
Here we briefly introduce the object functions of the four CG-SSL frameworks used in this paper, for a more detailed discussion about object functions including other graph contrastive learning and graph self-supervised learning frameworks which can refer to the survey papers [34, 33, 19]. We use the following notations:

### v2

@@ Appendix A Dataset and training configuration
Datasets. The node classification datasets used in this paper include the Cora, CiteSeer, and PubMed citation networks [16], as well as the Photo and Computers co-purchase networks [32]. Additionally, we use the Coauthor-CS and Coauthor-Phy co-author relationship networks. The statistics of node-level datasets are present in Table 4. The graph classification datasets include: The MUTAG dataset, which features seven types of graphs derived from 188 mutagenic compounds; the NCI1 dataset, which contains compounds tested for their ability to inhibit human tumor cell growth; the PROTEINS dataset, where nodes correspond to secondary structure elements connected if they are adjacent in 3D space; and the IMDB-BINARY and IMDB-MULTI movie collaboration datasets, where graphs depict interactions among actors and actresses, with edges denoting their collaborations in films. These movie graphs are labeled according to their genres. The statistics of graph-level datasets are present in Table 5. All datasets can be accessed through PyG library  [^1] https://pytorch-geometric.readthedocs.io/en/latest/modules/datasets.html. All experiments are conducted using 8 NVIDIA A100 GPU.
@@
Training configuration. For each CG-SSL framework, we implement it based on [48]   [^2] https://github.com/PyGCL/PyGCL. We use the following hyperparameters: the learning rate is set to $5\times 10^{-4}$, and the node hidden size is set to $512$, the number of GCN encoder layer is set $\in\{1,2\}$. For all node classification datasets, training epochs are set $\in\{50,100,150,200,400,1000\}$, and for all graph classification datasets, training epochs are set $\in\{20,40,...,200\}$. To achieve performance closer to the global optimum, we use randomized search to determine the optimal probability of edge perturbation and SPAN perturbation ratio. For Cora and CiteSeer  the search is conducted one hundred times, and for all other datasets, it is conducted twenty times. For all graph classification datasets, the batch size is set to $128$.
Appendix B Preliminaries of Graph Spectrum and SPAN
@@
$\max_{\bm{\mathcal{T}}_{1},\bm{\mathcal{T}}_{2}\in\mathcal{S}}\left\|\operatorname{eig}\left(\mathbf{L}_{1}\right)-\operatorname{eig}\left(\mathbf{L}_{2}\right)\right\|_{2}^{2}$ (8)
@@
Appendix C Objective function of GCL framework
Here we briefly introduce the objective functions of the four CG-SSL frameworks used in this paper, for a more detailed discussion about objective functions including other graph contrastive learning and graph self-supervised learning frameworks which can refer to the survey papers [40, 39, 22]. We use the following notations: